# Mini-InternVL: A Flexible-Transfer Pocket Multimodal Model with 5% Parameters and 90% Performance

### v1

@@ 4 Experiments
In this section, we present a comprehensive evaluation of our model’s multimodal understanding and reasoning capabilities across a variety of benchmarks. The benchmarks used in our study are categorized into four distinct types: OCR-related tasks, including DocVQA [74] OCRBench [109] and InfographicVQA [82]; chart and diagram understanding, including AI2D [80] and ChartQA [83]; general multimodal tasks, such as MMBench [107]; and multimodal reasoning, including MMMU [110] and MathVista [108]. Additionally, we report the average score of our model on the OpenCompass [116].
@@ 4.2.1 Multi-View Image-Based Autonomous Driving
Data Source Total Data Data Type
@@
LLaVA_Instruct_zh [^1] https://huggingface.co/datasets/openbmb/llava_zh 158k VQA
@@
In DriveLM-nuScenes, the images are captured from six different viewpoints. We effectively utilize dynamic resolution feature to accommodate this type of data. Specifically, Mini-InternVL supports splitting images into 448$\times$448-sized tiles based on their aspect ratio. As illustrated in Figure 3, we resize the image of each view to 896$\times$448 pixels and then combine these images in a fixed sequence, resulting in a final resolution of 2688$\times$896. This means that the images are automatically processed into 12 tiles, and an additional thumbnail is added to provide the model with global context. Furthermore, we mark the image of each view with text indicating its camera position, such as “CAM_FRON”.
As shown in Figure 3, DriveLM-nuScenes contains QA pairs with coordinates, thus we need to normalize them to a range of 0 to 1000 to align with the output of Mini-InternVL. In the dataset, objects are represented by c tags. We use a tailored prompt: “Objects are encoded using <c, CAM, [cx,cy]>, where c is the identifier, CAM indicates the camera where the object s center point is situated, and x, y represent the horizontal and vertical coordinates of the center point of the 2D bounding box.” to guide the model on the composition of c tags. The ground truth responses typically include bounding box annotations for questions requiring a list of all objects. Therefore, We annotate the bounding box as <box>[[x1,y1,y2,y3]]</box>, where <box> and </box> are special tokens in Mini-InternVL. Additionally, we incorporate general datasets to the training set, to prevent the model from losing its general domain perception capabilities, maintaining a 1:4 ratio of general to domain-specific data. The sources of the general datasets are shown in Table 3.
Finally, we conduct full-parameter fine-tuning of Mini-InternVL using 8 A100 GPUs, training the model for 1 epoch with a learning rate of 1e-5.
We report the performance of our model after transfer learning on the CVPR 2024 Autonomous Driving Challenge [105]. Furthermore, we evaluate our model on autonomous driving scenarios from MME-Realworld [124], where we separately assess its performance on Perception and Reasoning tasks.
@@ Results.
As shown in Table 5, in the autonomous driving scenarios of MME-Realworld, we observe that even when using only DriveLM as domain-specific training data, our model achieves an improvement of over 10 points. The transferred model surpass the best-performing model on this task, LLaVA-OneVision-7B [5], as well as several commercial closed-source models such as GPT-4o [4] and Claude 3.5 Sonnet [125], demonstrating the strong generalization capability of our model.
@@ 4.2.2 Autonomous Driving with Temporal Information
Using single-frame images alone is insufficient for accurate perception and prediction of vehicle behavior. Therefore, we explore temporal expansion. Specifically, we utilize instruction-following data constructed by DriveGPT4 [50] from the BDD-X dataset [127] as our training set, which comprises 26k video clips. Multiple video frames are organized as described in Section 3.3. Each training sample contains four aspects of question-answer data: Action Description, Action Justification, Speed Signal Prediction, and Turning Angle Signal Prediction. We set the proportion of general to domain-specific data at 1:1.
@@
Table 6: The results on action tasks of BDD-X dataset. We provide evaluation results on action description, action justification, and full-text generation(i.e.,combining description and justification). B4 stands for BLEU4. “DA” means model after domain adaptation on BDD-X.
@@ 4.2.3 Medical Image Question Answering
Table 8: The sources of our medical data. The table presents the data types and the sample sizes collected from each source. We sampled a total of 500k image-text pairs from multiple publicly available datasets of different data types as our medical training data.
@@
Table 9: The results of our model on GMAI-MMBench. The results of other models are taken from GMAI-MMBench Leaderboard. “DA” means model after domain adaptation on medical data. After supervised fine-tuning, our model shows significant improvement and outperforms several medical-specialized models (e.g., LLaVA-Med, RadFM) and some commercial closed-source models (e.g., Claude3-Opus) on most metrics.
@@ Results.
Table 11 presents the performance of our models on remote sensing VQA and visual grounding tasks. On the RSVQA task, our model demonstrates strong performance under both high-resolution and low-resolution conditions. Unlike existing remote sensing MLLMs such as GeoChat [43] and SkySenseGPT [144], which support only single-resolution images, our model leverages dynamic resolution to effectively benefit from high-resolution training data. Compared to traditional models in the remote sensing domain such as RSVQA [141], Bi-Modal [142], and EasyToHard [143] our model achieves superior scores on both RSVQA-HR-Test1 and RSVQA-HR-Test2, showcasing its generalization ability. Furthermore, our models of three different sizes outperform SkyEyeGPT on DIOR-RSVG, indicating that our framework can effectively model visual grounding tasks.
@@ 4.3 Ablation Study
Table 12: Comparison between Mini-InternVL-2B with Mini-InternVL-CLIP-2B across various tasks. We test the model on DriveLM val set after finetuning them using domain-specific data.
@@
Table 13: Comparing the per-GPU memory consumption and training speed of the 3 methods. Our experiment use ZeRO1 [145] and is conducted on 8 A100 GPUs.
@@
We demonstrate the effectiveness of knowledge distillation in this study. We train a new 2B-sized MLLM, which we refer to as Mini-InternVL-CLIP-2B. Specifically, we maintain the structure of the MLLM while replacing the visual encoder from InternVIT with CLIP-Large-336 [13], employing the same training recipe as described in Section 3.1. As shown in Table 12, the results across multiple benchmarks indicate that Mini-InternVL-CLIP-2B significantly outperforms Mini-InternVL-CLIP-2B on general benchmarks, particularly in document-related tasks. This clearly illustrates that our knowledge distillation in our method effectively enables InternVIT to acquire visual world knowledge. Furthermore, we adapt Mini-InternVL-CLIP-2B to autonomous driving tasks, and the results demonstrate the advantages of our model in proprietary domain transfer.
@@
In this study, we investigate the effect of the proportion between general data and domain-specific data on model transferability. We conducted experiments in the autonomous driving domain, utilizing all samples from DriveLM and supplementing them with general training data at multiples of $r$ times the number of DriveLM samples. The results are shown in Figure 4(a). Our findings indicate that relying exclusively on domain-specific training data does not yield optimal performance on downstream tasks. Introducing a specific ratio of general data not only enhances performance on domain-specific tasks but also reduces performance degradation on general multimodal benchmarks. This demonstrates that incorporating an appropriate proportion of general data is crucial for improving the model’s generalization ability and maintaining its general capabilities. For our autonomous driving scenario, we observe that performance peaks at $r=0.25$; beyond this point, performance slightly declines as $r$ increases. This suggests that we can achieve benefits without significantly increasing computational load.
@@
We investigate how varying the quantity of training samples affects performance on downstream tasks. In this experiment, we use different amounts of training data while maintaining a 1:1 ratio between general data and domain-specific data, as shown in Figure 4(b) Training the model with only one-quarter of the full dataset significantly reduces computational load during training while resulting in only a minor loss in performance. Notably, the model’s score on general benchmarks remains largely unchanged when the proportion of different training data is kept constant.
@@
In this study, we examine the effects of different adaptation methods LoRA, freezing the visual encoder, and full-parameter fine-tuning on model performance. Using the dataset described in  Section 4.2.4, we apply these three methods to train the model across varying numbers of steps and evaluate its performance on four benchmarks: general multimodal VQA, remote sensing VQA, and visual grounding. We record memory consumption and training speed during the process, as shown in  Table 13. As illustrated in  Figure 5, model performance converges after 1500 steps, with each method exhibiting distinct performance ceilings. Notably, full-parameter fine-tuning achieves the highest scores on domain-specific tasks. Additionally, we find that LoRA underperforms on the visual grounding task, and freezing the vision encoder strikes a balance between performance and computational efficiency. The scores for all three adaptation methods remain stable across different training steps on the general multimodal benchmark, maintaining strong performance in general domain even with extended training.
@@ 5 Conclusion
In this work, we introduce Mini-InternVL, a series of lightweight, open-source MLLMs designed to tackle the challenges of deploying MLLMs in resource-constrained environments. Mini-InternVL utilizes InternViT-300M as a compact visual encoder, integrating world knowledge across multiple domains through knowledge distillation from a more capable teacher model, thereby addressing the limitations of encoders like CLIP-ViT. Mini-InternVL achieves approximately 90% of the performance of larger models using significantly fewer parameters,excelling particularly in tasks such as OCR and domain-specific image understanding. To facilitate the application of small-scale multimodal models in specialized fields, we employe a unified transfer format, enabling our models to be effectively adapted to multiple specific domains, where they achieve comparable performance to other domain-specific approaches. We hope that this work provides valuable insights into the application of MLLM.

### v3

@@ 4 Experiments
In this section, we present a comprehensive evaluation of our model’s multimodal understanding and reasoning capabilities across a variety of benchmarks. The benchmarks used in our study are categorized into four distinct types: OCR-related tasks, including DocVQA [84], OCRBench [116] and InfographicVQA [92]; chart and diagram understanding, including AI2D [90] and ChartQA [30]; general multimodal tasks, such as MMBench [29]; and multimodal reasoning, including MMMU [117] and MathVista [31]. Additionally, we calculate the average score across these benchmarks, with the OCRBench score divided by 10.
@@ 4.2.1 Multi-View Image-Based Autonomous Driving
Data Source Size Data Type
@@
LLaVA-Instruct-ZH [^1] https://huggingface.co/datasets/openbmb/llava_zh 158K VQA
@@
In DriveLM-nuScenes, the images are captured from six different viewpoints. We effectively utilize dynamic resolution features to accommodate this type of data. Specifically, Mini-InternVL supports splitting images into 448$\times$448-sized tiles based on their aspect ratio. As illustrated in Figure 3, we resize the image of each view to 896$\times$448 pixels and then combine these images in a fixed sequence, resulting in a final resolution of 2688$\times$896. This means that the images are automatically processed into 12 tiles, and an additional thumbnail is added to provide the model with global context. Furthermore, we mark the image of each view with text indicating its camera position, such as “CAM_FRON”.
As shown in Figure 3, DriveLM-nuScenes contains QA pairs with coordinates, thus we need to normalize them to a range of 0 to 1000 to align with the output of Mini-InternVL. In the dataset, objects are represented by c tags. We use a tailored prompt: “Objects are encoded using <c, CAM, [cx,cy]>, where c is the identifier, CAM indicates the camera where the object’s center point is situated, and x, y represent the horizontal and vertical coordinates of the center point of the 2D bounding box.” to guide the model on the composition of c tags. The ground truth responses typically include bounding box annotations for questions requiring a list of all objects. Therefore, We annotate the bounding box as <box>[[x1,y1,y2,y3]]</box>, where <box> and </box> are special tokens in Mini-InternVL. Additionally, we incorporate general datasets into the training set, to prevent the model from losing its general domain perception capabilities, maintaining a 1:4 ratio of general to domain-specific data. The sources of the general datasets are shown in Table 3.
Finally, we conduct full-parameter fine-tuning of Mini-InternVL using 8 A100 GPUs, training the model for 1 epoch with a learning rate of 1e-5. We report the performance of our model after transfer learning on the CVPR 2024 Autonomous Driving Challenge [114]. Furthermore, we evaluate our model on autonomous driving scenarios from MME-Realworld [130], where we separately assess its performance on Perception and Reasoning tasks.
@@ Results.
As shown in Table 5, in the autonomous driving scenarios of MME-Realworld, we observe that even when using only DriveLM as domain-specific training data, our model achieves an improvement of over 10 points. The transferred model surpasses the best-performing model on this task, LLaVA-OneVision-7B [5], as well as several commercial closed-source models such as GPT-4o [4] and Claude 3.5 Sonnet [131], demonstrating the strong generalization capability of our model.
@@ 4.2.2 Autonomous Driving with Temporal Information
Using single-frame images alone is insufficient for accurate perception and prediction of vehicle behavior. Therefore, we explore temporal expansion. Specifically, we utilize instruction-following data constructed by DriveGPT4 [60] from the BDD-X dataset [133] as our training set, which comprises 26K video clips. Multiple video frames are organized as described in Section 3.3. Each training sample contains four aspects of question-answer data: Action Description, Action Justification, Speed Signal Prediction, and Turning Angle Signal Prediction. We set the proportion of general to domain-specific data at 2:1.
@@
Table 6: The results on action tasks of BDD-X dataset. We provide evaluation results on action description, action justification, and full-text generation (i.e., combining description and justification). “B4” stands for BLEU4. “DA” means model after domain adaptation on BDD-X.
@@ 4.2.3 Medical Image Question Answering
Table 8: The sources of our medical data. The table presents the data types and the sample sizes collected from each source. We sample a total of 500K image-text pairs from multiple publicly available datasets of different data types as our medical training data.
@@
Table 9: The results of our model on GMAI-MMBench. The results of other models are taken from the GMAI-MMBench leaderboard. “DA” means model after domain adaptation on medical data. After supervised fine-tuning, our model shows significant improvement and outperforms several medical-specialized models (e.g., LLaVA-Med, RadFM) and some commercial closed-source models (e.g., Claude3-Opus) on most metrics.
@@ Results.
Table 11 presents the performance of our models on remote sensing VQA and visual grounding tasks. On the RSVQA task, our model demonstrates strong performance under both high-resolution and low-resolution conditions. Unlike existing remote sensing MLLMs such as GeoChat [53] and SkySenseGPT [150], which support only single-resolution images, our model leverages dynamic resolution to effectively benefit from high-resolution training data. Compared to traditional models in the remote sensing domain–such as RSVQA [147], Bi-Modal [148], and EasyToHard [149]–our model achieves superior scores on both RSVQA-HR-Test1 and RSVQA-HR-Test2, showcasing its generalization ability. Furthermore, our models of three different sizes outperform SkyEyeGPT on DIOR-RSVG, indicating that our framework can effectively model visual grounding tasks.
@@ 4.3 Ablation Study
Table 12: Comparison between Mini-InternVL-2B with Mini-InternVL-CLIP-2B across various tasks. We test the model on DriveLM val set after fine-tuning them using domain-specific data.
@@
Table 13: Comparing the per-GPU memory usage and training speed of 3 methods. Our experiment use ZeRO1 [151] and is conducted on 8 A100 GPUs.
@@
We demonstrate the effectiveness of knowledge distillation in this study. We train a new 2B-sized MLLM, which we refer to as Mini-InternVL-CLIP-2B. Specifically, we maintain the structure of the MLLM while replacing the vision encoder from InternViT with CLIP-ViT-L-336px [18], employing the same training recipe as described in Section 3.1. As shown in Table 12, the results across multiple benchmarks indicate that Mini-InternVL-2B significantly outperforms Mini-InternVL-CLIP-2B on general benchmarks, particularly in document-related tasks. This clearly illustrates that our knowledge distillation in our method effectively enables InternViT to acquire visual world knowledge. Furthermore, we adapt both models to autonomous driving tasks, and the results demonstrate the advantages of our model in proprietary domain transfer.
@@
In this study, we investigate the effect of the proportion between general data and domain-specific data on model transferability. We conduct experiments in the autonomous driving domain, utilizing all samples from DriveLM and supplementing them with general training data at multiples of $r$ times the number of DriveLM samples. The results are shown in Figure 4(a). Our findings indicate that relying exclusively on domain-specific training data does not yield optimal performance on downstream tasks. Introducing a specific ratio of general data not only enhances performance on domain-specific tasks but also reduces performance degradation on general multimodal benchmarks. This demonstrates that incorporating an appropriate proportion of general data is crucial for improving the model’s generalization ability and maintaining its general capabilities. For our autonomous driving scenario, we observe that performance peaks at $r=0.25$; beyond this point, performance slightly declines as $r$ increases. This suggests that we can achieve benefits without significantly increasing computational load.
@@
We investigate how varying the quantity of training samples affects performance on downstream tasks. In this experiment, we use different amounts of training data while maintaining a 1:1 ratio between general data and domain-specific data, as shown in Figure 4(b). Training the model with only one-quarter of the full dataset significantly reduces the computational load during training while resulting in only a minor loss in performance. Notably, the model’s score on general benchmarks remains largely unchanged when the proportion of different training data is kept constant.
@@
In this study, we examine the effects of different adaptation methods–LoRA, freezing the vision encoder, and full-parameter fine-tuning–on model performance. Using the dataset described in Section 4.2.4, we apply these three methods to train the model across varying numbers of steps and evaluate its performance on three tasks: general multimodal VQA, remote sensing VQA, and visual grounding. We record memory consumption and training speed during the process, as shown in Table 13. As illustrated in Figure 5, model performance converges after 1500 steps, with each method exhibiting distinct performance ceilings. Notably, full-parameter fine-tuning achieves the highest scores on domain-specific tasks. Additionally, we find that LoRA underperforms on the visual grounding task, and freezing the vision encoder strikes a balance between performance and computational efficiency. The scores for all three adaptation methods remain stable across different training steps on the general multimodal benchmark, maintaining strong performance in the general domain even with extended training.
@@ 5 Conclusion
In this work, we introduce Mini-InternVL, a series of lightweight, open-source MLLMs designed to tackle the challenges of deploying MLLMs in resource-constrained environments. Mini-InternVL utilizes InternViT-300M as a compact vision encoder, integrating world knowledge across multiple domains through knowledge distillation from a more capable teacher model, thereby addressing the limitations of encoders like CLIP-ViT. Mini-InternVL achieves approximately 90% of the performance of larger models using significantly fewer parameters, excelling particularly in tasks such as OCR and domain-specific image understanding. To facilitate the application of small-scale multimodal models in specialized fields, we employ a unified transfer format, enabling our models to be effectively adapted to multiple specific domains, where they achieve comparable performance to other domain-specific approaches. We hope that this work provides valuable insights into the application of MLLM.